# Electroweak Monopole-Antimonopole Pair Production at LHC

### v2

@@ VII Discussion
An urgent task at LHC is to discover the electroweak monopole. Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it. But the problem is that the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV. And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it. In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition.
Our work in this paper tells that, if LHC produces the monopole by the topological thermal fluctuation of the Higgs vacuum, LHC could circumbent this energy constraint and produce the electroweak monopole even when the monopole mass becomes 11.0 TeV. This is because in this thermal production the monopole mass depends on the temperature at production. In this case the mass of the monopoles produced right after the electroweak phase transition (the baby monopole mass) could be as low as 1.0 TeV, even when the monopole mass at zero temperature (the adolescent monopole mass) becomes 11.0 TeV.
So, if the monopole production mechanism at LHC is the thermal fluctuation of the Higgs vacuum (the same mechanism as the monopole production in the early universe), the present 14 TeV LHC could produce the electroweak monopole, and thus MoEDAL could detect it. Our results suggest that LHC could produce the infant monopoles of mass around 5.3 TeV in average (assuming the adolescent monopole mass is 11 TeV). This is really remarkable, because this implies that there is a finite possibility that the MoEDAL detector at LHC could actually detect the infant electroweak monopoles, even though it is not likely to detect the adolescent (i.e., grown-up) monopoles.
It should be emphasized that this is a unique feature of the finite temperature physics which is absent in the zero temperature physics. In particular, this is impossible if the monopole production mechanism at LHC is the Drell-Yan or Schwinger mechanism, where only the adolescent monopole enters. From this we can definitely say that the detection of the infant monopole at LHC should be interpreted as the confirmation that the electroweak monopole production mechanism is indeed topological.
Another important result of our work is that there is a real possibility for LHC to produce the monopoles in the form of the monopolia. When LHC produces the monopoles, it should produce them in pairs. But since they have strong magnetic attraction some of them could form bound states, although many of the monopole-antimonopole pairs could quickly annihilate. An interesting feature of this monopolium is that the mass could be considerably lighter than the mass of the monopole-antimonopole pair, because the monopolium binding energy could be of the order of ten TeV. In fact, our study in this paper suggests that the binding energy of the ground state could be as much as 16.7 TeV, so that the monopole-antimonopole pair of the mass 22.0 TeV could form a bound state of the mass 5.3 TeV. This provides another way for LHC to circumbent the energy constraint and produce the monopole pairs.
This tells that there are two ways for LHC to produce the monopole pairs, the monopole and antimonopole as independent particles or the monopole-antimonopole pair together as a monopolium bound state. So MoEDAL and ATLAS could try to detect the monopole (and antimonopole) separately, or together as a monopolium.
Detecting the monopole (and antimonopole) should be straightforward in principle, but detecting the monopolium could be tricky. This is because the monopolium has zero magnetic charge, so that we have to look for the magnetically neutral particles of mass of few TeV. Moreover, the monopole and antimonopole inside the monopolium could easily annihilate to photons and electron-positron pairs. In this case we may have to look for the decay modes of the monopolium.
One might wonder if we could also have the infant monopoles in cosmology produced in the early universe. This may not be possible, because the monopoles in the early universe are in contact with the hot thermal radiation bath which feeds up the energy and allows them to grow up. This means that the remnant monopoles from the early universe that we could detect at present are expected to be the adolescent monopoles. This implies that the quenched infant monopole at LHC is not a generic phenomenon but an outcome of the limited energy and size of the p-p fireball at LHC.
Admittedly the idea of the infant monopole at LHC sounds strange and appears hard to swallow. We emphasize, however, that this idea actually comes from the fact that in the electroweak phase transition the W (and Higg) boson mass becomes temperature dependent. Once we accept this, the idea of the infant monopole follows naturally. Moreover, it appears that (as far as we know) logically it is difficult to exclude this possibility. Ultimately, of course, only experiment could tell whether this idea is realistic or not.
We hope that our results in this paper could help MoEDAL to detect the electroweak monopole at LHC.
@@ ACKNOWLEDGEMENT
PB and FB are supported by the Institute of Experimental and Applied Physics, Czech Technical University in Prague and the Research Centre for Theoretical Physics and Astrophysics, Institute of Physics, Silesian University in Opava. YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University, Korea. The authors thank their supports.

### v4

@@ VII Discussion
An urgent task at LHC is to discover the electroweak monopole. Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it. But the problem is that intuitively the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV. And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it. In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition [15, 18].
In this paper we have shown that the question if the present LHC could produce the monopole or not critically depends on the monopole production mechanism at LHC. Our work in this paper tells the followings. First, if the monopole production mechanism at LHC is the topological thermal fluctuation of the Higgs vacuum, there is practically no hope that it could produce the monopole, simply because the temperature of the p-p fireball is too low to produce the monopole thermally. This is unexpected, but if the recent ALICE measurement of the temperature of the p-p fireball is trustable, this conclusion is unavoidable.
Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole if the monopole mass becomes less than 7 TeV. This was expected. What is new is that, more probably LHC could produce the monopole in the form of the monopolium bound states for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV. This is because the monopolium binding energy greatly reduces the mass of the monopolium.
This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole even when the monopole mass becomes bigger than 7 TeV. This is remarkable.
Of course, with the Schwinger mechanism at LHC we can have the same conclusion, if the p-p fireball could create strong enough magnetic background which can actually produce the monopole-antimonopole pairs. But it is by no means clear how the p-p fireball at LHC could generate such a strong magnetic background.
To summarize, our results tells that a most probable way for LHC to produce the monopole is the Drell-Yan process, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states than as individual monopoles. In this case the most realistic way to detect the monopole at LHC is to detect the two photon decay modes of the monopoliums. And in principle ATLAS and CMS could do that.
But there is the possibility that the LHC (with the next energy upgrading) may not produce the monopole signal at all. This could mean that the monopole production mechanism at LHC is probably not the Drell-Yan process. In this case the only way to detect the electroweak monopole could be to look for the remnant monopoles produced during the early universe, with the “cosmic” MoEDAL. This makes the detection of the remnant monopoles produced during the early universe very important.
@@ ACKNOWLEDGEMENT
PB and FB are supported by the Institute of Experimental and Applied Physics, Czech Technical University in Prague and the Research Centre for Theoretical Physics and Astrophysics, Institute of Physics, Silesian University in Opava. YMC is supported in part by the National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University, Korea.